OpenStudy (anonymous):

Solve for x −6 x 18= 4(x+8)

OpenStudy (anonymous):

I got to -6x + 18= 4x + 32 and can't figure out what to do next

OpenStudy (butterflydreamer):

From here, you want to make "x" the subject or isolate x. :) \[-6x + 18 = 4x + 32\] To do this, you would ADD 6x to both sides \[-6x + 18 + 6x = 4x + 32 + 6x\] Collect like terms: \[18 = 10x + 32\] Now, subtract 32 from both sides to make "10x" the subject. \[18 - 32 = 10x + 32 - 32\] \[18 - 32 = 10x\] Do you know where to go from here? ^_^

OpenStudy (anonymous):

Yes thank you